OpenStudy (anonymous):

Combine the like terms to make a simpler expression: -2q+q

OpenStudy (anonymous):

q(1-2) is the same as -2q+2

OpenStudy (midhun.madhu1987):

-2+1=?

OpenStudy (anonymous):

so the answer is q(1-2)

OpenStudy (anonymous):

thank you @midhun.madhu1987

OpenStudy (anonymous):

so -2q+1 would equal 3

OpenStudy (anonymous):

-2q+q = -q

OpenStudy (midhun.madhu1987):

-2q+q means -2q+1q but -2+1 = -1 so -2q+1q = ?

OpenStudy (anonymous):

thank you @midhun.madhu1987

OpenStudy (midhun.madhu1987):

yw :)